OpenStudy (gabylovesyou):

Choose the slope-intercept equation of the line that passes through the point (-2, 4) and is parallel to y = -3x + 6. y = 1/3x + 14/3 y = 3x + 10 y = -3x - 2 y = -1/3x + 10/3

OpenStudy (anonymous):

Parallel lines have the same slope. Put the information you know into point-slope form: y-k=m(x-h) m is the slope the line passes through the point (h,k) y-4=-3(x-(-2)) y-4=-3x-6 y=-3x-2

OpenStudy (anonymous):

The first concept we need for this is slope intercept of a line. y = mx+b where m is the slope, and b is the y-intercept. The second concept we need for this is the fact that parallel lines have the same slope.

OpenStudy (anonymous):

74

OpenStudy (gabylovesyou):

WTH @MYnameisBrandon ?? lol

OpenStudy (gabylovesyou):

Thank u guys..... kisses xoxoxo muaaa

OpenStudy (anonymous):

it is!!!!!!!!!!!!!!

OpenStudy (gabylovesyou):

no it isnt stupid...

OpenStudy (gabylovesyou):

so its C :)

OpenStudy (gabylovesyou):

right?

OpenStudy (anonymous):

@Copythat it's generally better to lead a student to an answer than it is to directly give them the answer. Maybe in the future give some explanation and then see if they can find the answer on their own.